# Austria’s KlimaTicket: Assessing the short-term impact of a cheap nationwide travel pass on demand

### 12 Jan 2024 v1

@@
Abstract: Measures to reduce transport-related greenhouse gas emissions are of great importance to policy-makers. A recent example is the nationwide KlimaTicket in Austria, a country with a relatively high share of transport-related emissions. The cheap yearly season ticket introduced in October 2021 allows unlimited access to Austria’s public transport network. Using the synthetic control and synthetic difference-in-differences methods, I assess the causal effect of this policy on public transport demand by constructing a data-driven counterfactual out of European railway companies to mimic the number of passengers of the Austrian Federal Railways without the KlimaTicket. The results indicate public transport demand grew slightly faster in Austria, i.e., 3.3 or 6.8 percentage points, depending on the method, than it would have in the absence of the KlimaTicket. However, the growth effect after the COVID-19 pandemic appears only statistically significant when applying the synthetic control method, and the positive effect on public transport demand growth disappears in 2022. Keywords: Public transport; Policy evaluation; Synthetic control method; Case study; Natural experiment Acknowledgments: Michael Stöckli provided excellent research assistance. I would like to thank Kevin Blättler, Silvio Sticher, and Widar von Arx for interesting discussions and helpful comments.
@@ 1 Introduction
While the effects of the KlimaTicket have already been discussed in first studies (see, e.g., Follmer and Justin, 2023, OECD IPAC International Programme for Action on Climate, 2022), I am the first to estimate the causal effect of the ticket on public transport demand. In this study, I show initial causal demand effects and how future analyses by researchers and practitioners can be structured to analyze the KlimaTicket or public transport price policies in general. In a comparative case study, I use the synthetic control and difference-in-differences methods, introduced by Abadie, Diamond, and Hainmueller (2010) or Arkhangelsky, Athey, Hirshberg, Imbens, and Wager (2019), to construct a synthetic Austrian Federal Railways (ÖBB). This synthetic counterfactual—a weighted subset of comparable European public transport companies—mimics the demand the actual ÖBB would have experienced in the absence of the KlimaTicket. Comparing the outcomes of the actual and the synthetic ÖBB, I estimate that the public transport demand in Austria grew on average 6.3 or 3.3 percentage points (depending on the method) faster after the COVID-19 pandemic, mainly driven by higher growth rates in 2021. However, I also show that statistical significance is questionable. To sum up, the causal effect of the KlimaTicket on public transport is positive but small and, from a statistical point of view, has to be taken with a grain of salt.
@@ 5 Discussion and Conclusion
In this study, I assessed the effect of the KlimaTicket on demand during the first and second years, right at the beginning and after the introduction. Therefore, I applied the synthetic control and synthetic difference-in-differences methods to mimic the ÖBBs’ public transport demand growth rates in the absence of the KlimaTicket. The effect of the KlimaTicket on the COVID-19 pandemic recovery is, from an economic perspective, positive but small. Statistically speaking, considering the high standard errors (mainly when applying the synthetic difference-in-differences method), chance might drive the effect. Moreover, in 2022, the positive causal effect on demand growth disappears (and even becomes slightly negative).
@@
A limitation of this study is the data availability. Only data from five companies was available to create the synthetic counterfactual. However, by looking at the pre-treatment period in Figure 2, I conclude that I was able to create an accurate approximation of the ÖBB. Nonetheless, I encourage future research projects with more units in the donor pool (I am also open to feedback regarding the revision of the underlying first draft of the paper). Moreover, I do not account (because of data availability reasons) for supply changes, e.g., frequency increases. Therefore, future studies should also account for supply changes in order to block off changing frequencies as an alternate explanation of the causal effect (see, e.g., Wallimann, Blättler, and von Arx, 2023, using an aggregate metric that inherits changes in public transport supply).
@@
Finally, I want to add some thoughts to the design of the KlimaTicket, i.e., a season ticket with a single fixed fee regardless of usage—without discussing revenues of the ÖBB and following implications on public transport subsidies being a topic for future research. From a pricing psychological point of view, such an offer can make sense for several reasons (see, e.g. Wirtz, Vortisch, and Chlond, 2015, Dang, Steffen, Weibel, and von Arx, 2024). For instance, no so-called taximeter effect exists when a customer possesses the KlimaTicket, i.e., she does not have to travel with a pay-per-use tariff, making the costs of each journey transparent and thus less attractive. Moreover, purchasing the KlimaTicket makes budget planning for a customer easier. Thus, I can talk about the so-called insurance effect, i.e., customers buy the KlimaTicket because they value mentally the agony of overspending more than the excitement of saving money (through pay-per-use traveling).
@@ Appendix A Contacted companies
Appendix B Robustness Investigation

### 09 Feb 2024 v2

@@
Abstract: Measures to reduce transport-related greenhouse gas emissions are of great importance to policy-makers. A recent example is the nationwide KlimaTicket in Austria, a country with a relatively high share of transport-related emissions. The cheap yearly season ticket introduced in October 2021 allows unlimited access to Austria’s public transport network. Using the synthetic control and synthetic difference-in-differences methods, I assess the causal effect of this policy on public transport demand by constructing a data-driven counterfactual out of European railway companies to mimic the number of passengers of the Austrian Federal Railways without the KlimaTicket. The results indicate public transport demand grew slightly faster in Austria, i.e., 3.3 or 6.8 percentage points, depending on the method, than it would have in the absence of the KlimaTicket. However, the growth effect after the COVID-19 pandemic appears only statistically significant when applying the synthetic control method, and the positive effect on public transport demand growth disappears in 2022.
Keywords: Public transport; Policy evaluation; Synthetic control method; Case study; Natural experiment Acknowledgments: Michael Stöckli provided excellent research assistance. I would like to thank Kevin Blättler, Silvio Sticher, and Widar von Arx for interesting discussions and helpful comments.
@@ 1 Introduction
While the effects of the KlimaTicket have already been discussed in first studies (see, e.g., Follmer and Justin, 2023, OECD IPAC International Programme for Action on Climate, 2022), I am the first to estimate the causal effect of the ticket on public transport demand. In this study, I show initial causal demand effects and how future analyses by researchers and practitioners can be structured to analyze the KlimaTicket or public transport price policies in general. In a comparative case study, I use the synthetic control and difference-in-differences methods, introduced by Abadie, Diamond, and Hainmueller (2010) or Arkhangelsky, Athey, Hirshberg, Imbens, and Wager (2019), to construct a synthetic Austrian Federal Railways (ÖBB). This synthetic counterfactual—a weighted subset of comparable European public transport companies—mimics the demand the actual ÖBB would have experienced in the absence of the KlimaTicket. Comparing the outcomes of the actual and the synthetic ÖBB, I estimate that the public transport demand in Austria grew on average 6.3 or 3.3 percentage points (depending on the method) faster after the COVID-19 pandemic, mainly driven by higher growth rates in 2021. However, I also show that statistical significance is questionable, inter alia by replacing the outcome variable with Eurostat data on rail passengers from many European countries. To sum up, the causal effect of the KlimaTicket on public transport is positive but small and, from a statistical point of view, has to be taken with a grain of salt.
@@ 3 Data
Finally, in order to challenge the data source, I re-estimate the effect by replacing the outcome variable with the rail passenger data provided by Eurostat. [^2] See, e.g., , accessed on January 17, 2024. The thing to notice is that I now observe 19 control units (i.e., all European countries with more than 10 million passengers per year from 2015 to 2022 and non-confidential data), including Bulgaria, Croatia, Czechia, Denmark, Finland, France, Germany, Greece, Ireland, Italy, Latvia, Luxembourg, Norway, Portugal, Romania, Slovakia, Spain, Sweden, and Switzerland.
@@ 4.2 Robustness investigations
As a second robustness check, I substitute the number of passengers from annual reports with the national rail passengers per country (see also Figure B.2 in Appendix B). Interestingly, the effect of the KlimaTicket on demand growth amounts to -3.8 and 0.2 for the synthetic control and synthetic difference-in-differences method and is, therefore, even negative once. With again high standard errors of 14.4 and 11.4, when applying the Eurostat data, I conclude that we see no effect of the KlimaTicket on rail passengers.
@@ 5 Discussion and Conclusion
In this study, I assessed the effect of the KlimaTicket on demand during the first and second years, right at the beginning and after the introduction. Therefore, I applied the synthetic control and synthetic difference-in-differences methods to mimic the ÖBBs’ public transport demand growth rates in the absence of the KlimaTicket. The effect of the KlimaTicket on the COVID-19 pandemic recovery is, from an economic perspective, positive but small. Statistically speaking, considering the high standard errors (mainly when applying the synthetic difference-in-differences method), chance might drive the effect. Moreover, in 2022, the positive causal effect on demand growth disappears (and even becomes slightly negative). Finally, I have shown in a robustness investigation that there is no (positive) empirical effect of the KlimaTicket on passenger numbers if we look at the national level of railway travelers.
@@
A limitation of this study is the data availability. Only data from five companies was available to create the synthetic counterfactual. However, by looking at the pre-treatment period in Figure 2, I conclude that I was able to create an accurate approximation of the ÖBB. In addition, regarding the data of the robustness check, I was also able to create a synthetic Austria that mimics the rail passenger growth in Austria. Moreover, I do not account (because of data availability reasons) for supply changes, e.g., frequency increases. Therefore, future studies should also account for supply changes in order to block off changing frequencies as an alternate explanation of the causal effect (see, e.g., Wallimann, Blättler, and von Arx, 2023, using an aggregate metric that inherits changes in public transport supply).
@@
Finally, I want to add some thoughts to the design of the KlimaTicket, i.e., a season ticket with a single fixed fee regardless of usage—without discussing revenues of the ÖBB and following implications on public transport subsidies being a topic for future research. From a pricing psychological point of view, such an offer can make sense for several reasons (see, e.g. Wirtz, Vortisch, and Chlond, 2015, Dang, Steffen, Weibel, and von Arx, 2024). For instance, no so-called taximeter effect exists when a customer possesses the KlimaTicket, i.e., she does not have to travel with a pay-per-use tariff, making the costs of each journey transparent and thus less attractive. Moreover, purchasing the KlimaTicket makes budget planning for a customer easier. This links to the insurance effect, summarizing the finding that customers buy the KlimaTicket because they mentally value the agony of overspending more than the excitement of saving money (through pay-per-use traveling).
@@ Appendix A Contacted companies
Appendix B Robustness investigations
@@
Figure B.2: Rail demand development of Austria and the synthetic Austria (Eurostat data)